# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

### v2

@@ Appendix F Backbone Methods & Experimental Details in Any-Shot Learning
F.1 Model Agnostic Meta Learning
In sinusoid regression, Meta-RL, MAML is used as the backbone algorithm. As previously discussed, MAML is widely applied in solving few-shot learning tasks. In mathematics, its optimization objective can be characterized as:
@@ F.1 Model Agnostic Meta Learning
When it comes to reinforcement learning scenarios, $\mathcal{D}_{\tau}$ corresponds to episodic returns collected from MDPs with either the meta policy or the fast adapted policy.
@@ F.3 Multi-Modal Prompt Learning
The many-shot prompt learning setting refers to the 16-shot classification task proposed by 96, 97. Based on the MaPLe backbone 44, we again freeze the CLIP model parameters and tune multi-modal prompts. The prompts are optimized on a selected ImageNet subset, with 16 samples from each category.
@@ G.4 Robotic DR
The machine learner is a neural network with 2 hidden layers of size 10 with the Rectified Linear Unit (ReLU) nonlinear activation units. The task batch for ERM, and GDRM is 10, and that for DRM is 20 as default. The temperature parameter in GDRM is 0.01. The learning rates for actor and critic are 3e-4. The following is about extra optimization details or setups in MPTS. The task identifier is encoded into the latent variable $\bm{z}\in\mathbb{R}^{x}$. The batch size of the identifier in training is 25 for LunarLander and 250 for ErgoReacher. The Lagrange multiplier is set as $\beta=1.0$, and we use the Adam optimizer with the learning rate 0.005 to update the risk learner.
@@ Appendix I Competing Interests & Author Contributions
Under the supervision of professor X.J., Q.C.W. conceptualized the idea of MPTS, designed the computational framework, formulated the mathematical part, and wrote the draft. Z.X. and J.S. implemented MPTS in sinusoid regression, prompt-based few-shot image classification, and PEFT image classification, collected experimental results to visualize, and added implementation details in Supplementary Material. Y.M., Y.Q., and Y.L. implemented MPTS in sinusoid regression, Meta-RL and robotic DR, collected experimental results to visualize, and added implementation details in Supplementary Material. X.J. supervised the progress of MPTS, organized technical discussions, reviewed and edited the original draft. All authors have read the manuscript and approved the public version.
@@
Xiangyang Ji received the B.S. degree in materials science and the M.S. degree in computer science from the Harbin Institute of Technology, Harbin, China, in 1999 and 2001, respectively, and the Ph.D. degree in computer science from the Institute of Computing Technology, Chinese Academy of Sciences, in 2008. He joined Tsinghua University, Beijing, in 2008, where he is currently a Professor with the Department of Automation, School of Informtion Science and Technology. He has authored more than 100 refereed conference and journal papers. His current research interests include signal processing, computer vision, computational photography, and intelligent decision-making.

### v3

@@ Appendix F Backbone Methods & Experimental Details in Any-Shot Learning
F.1 MAML
In sinusoid regression and Meta-RL, MAML is used as the backbone algorithm. As previously discussed, MAML is widely applied in solving few-shot learning tasks. In mathematics, its optimization objective can be characterized as:
@@ F.1 MAML
When it comes to reinforcement learning scenarios, $\mathcal{D}_{\tau}$ corresponds to episodic returns collected from MDPs with either the meta policy or the fast adapted policy. To ensure enough coverage of task space, we adopt a mixture strategy of MPTS and random sampling as an empirical regularizer in all RL scenarios, which is similar to work 33.
@@ F.3 Multi-Modal Prompt Learning
The many-shot prompt learning setting refers to the 16-shot classification task proposed in work 96, 97. Based on the MaPLe backbone 44, we again freeze the CLIP model parameters and tune multi-modal prompts. The prompts are optimized on a selected ImageNet subset, with 16 samples from each category.
@@ Appendix G Experimental Setups & Implementation Details
Visualized Results during Training Phases.
Note that the active selection and the random sampler with different batches affect the reflection of the machine learner’s performance. Hence, learning curves in sinusoid regression (Fig. 3.a), Meta-RL (Fig. 5.a-b), and DR (Fig. 6.a-b) are actually evaluated in a uniformly sampled validation task dataset for fair comparison. Details on these validation task dataset are attached in the opensourced code.
@@ G.4 Robotic DR
The machine learner is a neural network with 2 hidden layers of size 10 with the Rectified Linear Unit (ReLU) nonlinear activation units. The task batch for ERM, and GDRM is 10, and that for DRM is 20 as default. The temperature parameter in GDRM is 0.01. The learning rates for actor and critic are 3e-4. The following is about extra optimization details or setups in MPTS. The task identifier is encoded into the latent variable $\bm{z}$. The batch size of the identifier in training is 25 for LunarLander and 250 for ErgoReacher. The Lagrange multiplier is set as $\beta=1.0$, and we use the Adam optimizer with the learning rate 0.005 to update the risk learner.
@@ Appendix I Competing Interests & Author Contributions
Under the supervision of professor X.J., Q.C.W. conceptualized the idea of MPTS, designed the computational framework, formulated the mathematical part, and wrote the draft. Z.X. and J.S. implemented MPTS in sinusoid regression, prompt-based few-shot image classification, and PEFT image classification, collected experimental results to visualize, and added implementation details in Supplementary Material. Y.M., Y.Q., and Y.L. implemented MPTS in sinusoid regression, Meta-RL and robotic DR, collected experimental results to visualize, and added implementation details in Supplementary Material. X.J. supervised the progress of MPTS, organized technical discussions, reviewed and revised the original draft. All authors have read the manuscript and approved the public version.
@@
Xiangyang Ji received the B.S. degree in materials science and the M.S. degree in computer science from the Harbin Institute of Technology, Harbin, China, in 1999 and 2001, respectively, and the Ph.D. degree in computer science from the Institute of Computing Technology, Chinese Academy of Sciences, in 2008. He joined Tsinghua University, Beijing, in 2008, where he is currently a Professor with the Department of Automation, School of Information Science and Technology. He has authored more than 100 refereed conference and journal papers. His current research interests include signal processing, computer vision, computational photography, and intelligent decision-making.